OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 3 sin πt + 4 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) (a) Find the average velocity during each time period. (i) [1, 2]

OpenStudy (anonymous):

\[v(s)=s \prime(t)\] average velocity is the gradient \[f(x)-f(x _{1})/x-x _{1}=f(1)-f(2)/(1-2)\] \[s(1)=3\sin \pi+4\cos \pi=-4\] \[s(2)=3\sin 2\pi+4\cos 2\pi=4\] \[s(1)-s(2)/1-2=-4-4/-1=8\] 8cm per second